OpenStudy (anonymous):

find critical points of (x+1)/(x^(2)+1)

OpenStudy (anonymous):

i got the derivative as [(x^(2)+1)-(-2x^(2)+2x)]/[(x^(2)+1]^2

OpenStudy (anonymous):

Then you need to take that and make it equal 0. Find just when the top = 0. Then look for the bottom and if any of the answers match then they are holes. (x+3)(x-5) __________ =0 (x-5) x=-3 and 5. But if x =5 then the bottom = 0 and therefore that isn't answer.

OpenStudy (anonymous):

so you first solve numerator=0 and then denominator=0? and for this one there are no critical points right?

OpenStudy (anonymous):

@malical

OpenStudy (anonymous):

First off it looks like your derivative is wrong. Denominator(derivative numerator)-(numerator)(derivative denominator) over denominator squared. If you did do this correctly I'm sorry. I didn't actually work it out, it just smelled wrong to me.